OpenStudy (anonymous):

2sqaure root6 /2squre root2 - 2sqaure root2/2sqaure root 2 =?

OpenStudy (anonymous):

let me clarify the question.... is it like this: \(\large \frac{2\sqrt6}{2\sqrt2}-\frac{2\sqrt2}{2\sqrt2} \) ????

OpenStudy (anonymous):

Just combine them..

OpenStudy (anonymous):

As the denominator is same, just combine them.. And subtract the numerator..

OpenStudy (anonymous):

yh i did but with the denominator i always get confused with removing the root signs

OpenStudy (anonymous):

Don't worry, talented @dpaInc is here to help you..

OpenStudy (anonymous):

yes... denominators are the same, so just subtract the numerator... factor out a 2 from the numerator and cancel with the 2 in the denominator... but not as talented as @waterineyes , !!!

OpenStudy (anonymous):

awww... what happened? i was waiting for a response....

OpenStudy (anonymous):

\[2(\sqrt{6} - \sqrt{2}) / 2\sqrt{2} - \sqrt{2})\]

OpenStudy (anonymous):

like that ?

OpenStudy (anonymous):

waterineyes is talented?? Really? Breaking News for today..

OpenStudy (anonymous):

the denominator is incorrect.... here... \(\large \frac{2\sqrt6}{2\sqrt2}-\frac{2\sqrt2}{2\sqrt2}=\frac{2\sqrt6-2\sqrt2}{2\sqrt2}=\frac{2(\sqrt6-\sqrt2)}{2\sqrt2}=\frac{\cancel2^1(\sqrt6-\sqrt2)}{\cancel2^1\sqrt2}=\frac{\sqrt6-\sqrt2}{\sqrt2} \) now we must rationalize the denominator...

OpenStudy (anonymous):

and yes... talented.... it's old news.... :)

OpenStudy (anonymous):

how do we know when to rationalize the denominator ?

OpenStudy (anonymous):

simplest radical form means there are no radical(s) in the denominator of a rational expression.

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

See if you have radical in the denominator, then to make the denominator you will rationalize it?? Getting??

OpenStudy (anonymous):

yeah okay :)

OpenStudy (anonymous):

For example: \[\frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

Here you can rationalize it...

OpenStudy (anonymous):

so then do we get sqaure root 12- sqaure root 4 over 4

OpenStudy (anonymous):

hmmm .... now that i think about it, we went about this the long way when we could have just done it from the beginning....

OpenStudy (anonymous):

Can't we take sqrt{2} common from numerator? @dpaInc

OpenStudy (anonymous):

when you multiply the denominator square root 2 by square root 2 do u just get 2 ? or 2x2

OpenStudy (anonymous):

\[\frac{2\sqrt6}{2\sqrt2}-\frac{2\sqrt2}{2\sqrt2}=\frac{2\sqrt6-2\sqrt2}{2\sqrt2}=\frac{2(\sqrt6-\sqrt2)}{2\sqrt2}=\frac{\cancel2^1(\sqrt6-\sqrt2)}{\cancel2^1\sqrt2}=\frac{\sqrt6-\sqrt2}{\sqrt2}\]

OpenStudy (anonymous):

\[\frac{\sqrt6-\sqrt2}{\sqrt2} = \frac{\sqrt{2}(\sqrt{3} - 1)}{\sqrt{2}}\]

OpenStudy (anonymous):

like this... \(\large \frac{2\sqrt6}{2\sqrt2}-\frac{2\sqrt2}{2\sqrt2}=\frac{\cancel2\sqrt6}{\cancel2\sqrt2}-1=\frac{\cancel{\sqrt2} \sqrt3}{\cancel{\sqrt2}}-1=\sqrt3 - 1 \)

OpenStudy (anonymous):

\[\sqrt{6} = \sqrt{2 \times 3} = \sqrt{2} \cdot \sqrt{3}\]

OpenStudy (anonymous):

haha.... many ways to do it... what I tell u??? the talented @waterineyes .... :)

OpenStudy (anonymous):

im really confused now :(

OpenStudy (anonymous):

can you please show me one step in order

OpenStudy (anonymous):

ok... i think the shortest and fastest way is that last way i showed.... \(\large \frac{2\sqrt6}{2\sqrt2}-\frac{2\sqrt2}{2\sqrt2}=\frac{\cancel2\sqrt6}{\cancel2\sqrt2}-1=\frac{\cancel{\sqrt2} \sqrt3}{\cancel{\sqrt2}}-1=\sqrt3 - 1 \) do you undestand the flow?

OpenStudy (anonymous):

how did u get -1 ?

OpenStudy (anonymous):

ohh yhhh i get it

OpenStudy (anonymous):

\(\large \frac{2\sqrt2}{2\sqrt2}=1 \) dooh... you already rplied...

OpenStudy (anonymous):

the answer has square root 6 - square root 4 over 4 ?

OpenStudy (anonymous):

this is the answer: \(\large \frac{\sqrt6 - \sqrt4}{4} \) ????

OpenStudy (anonymous):

yeah its in the text book ?

OpenStudy (anonymous):

you sure? i don't think thats right...

OpenStudy (anonymous):

lemme double check....

OpenStudy (anonymous):

thats what it says

OpenStudy (anonymous):

Do you have scanner?? Can you here upload the copy of question ??

OpenStudy (anonymous):

If we rationalize it then also we get: \[\frac{2 \sqrt{3} - 2}{2}\]

OpenStudy (anonymous):

its questions 3b

OpenStudy (anonymous):

that's a trig expression... not the radical thingy you gave...

OpenStudy (anonymous):

yeah if u solve it then u sure get what i gave u

OpenStudy (anonymous):

should *

OpenStudy (anonymous):

not quite... the first product is sqrt3/2 * sqrt2/2 = sqrt6/4

OpenStudy (anonymous):

i got square root 3 over2 x 1pver square root2 - 1/2 x 1/square root 2

OpenStudy (anonymous):

ok.. everything is messed can we start with the original question from the book?

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

ok... here's question 3b: \(\large sin(\pi/3)cos(\pi/4)-cos(\pi/3)sin(\pi/4) \) = \(\large (\frac{\sqrt3}{2})(\frac{\sqrt2}{2})-(\frac{1}{2})(\frac{\sqrt2}{2}) \) is this ok so far?

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

wait sorry

OpenStudy (anonymous):

for sin pi/4 i have 1/sqaure root 2

OpenStudy (anonymous):

\(\large \frac{1}{\sqrt2}=\frac{\sqrt2}{2} \) they're the same value only the right side is in simplest radical form

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

\(\large sin(\pi/3)cos(\pi/4)-cos(\pi/3)sin(\pi/4) \) = \(\large (\frac{\sqrt3}{2})(\frac{\sqrt2}{2})-(\frac{1}{2})(\frac{\sqrt2}{2}) \) = \(\large \frac{\sqrt6}{4}-\frac{\sqrt2}{4} \) = \(\large \frac{\sqrt6-\sqrt2}{4}\)

OpenStudy (anonymous):

thank uu

OpenStudy (anonymous):

and that answer matches what wolfram alpha has... but it still does not match your answerbook... i'm confident that this answer is correct... yw.... :)

OpenStudy (anonymous):

no thats the correct answer

OpenStudy (anonymous):

ok...:)